OpenStudy (anonymous):

Joe has 4 red pens, 1 green pen, 2 blue pens, and 3 black pens. What is the probability he will randomly select a black pen and then,without replacing it, randomly selecti a blue pen?

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

OH I'M SORRY. I'M NOT GOOD WITH PROBABILITY

OpenStudy (anonymous):

ITS OK ARI

OpenStudy (raden):

P(Black) = 3/10 P(Blue) = 2/9 multiply of both

OpenStudy (anonymous):

event A: selecting "1" black from a total of (4+1+2+3)=10; \[P(A)={3\over10}\] event B: selecting "1" blue from the remainning "9" \[P(B)={2\over9}\] now, the events are mutually independent... \[P(A\cap B)=P(A)\times P(B)={3\over10}\times{2\over9}\]

OpenStudy (anonymous):

Oh, ok, thanks!

OpenStudy (anonymous):

IDK who to give BEST ANSWER to. LOL.

OpenStudy (anonymous):